2X + 6Y = 0
6X - 2Y = 0
I DO NOT UNDERSTAND. I HAVE TO GRAPH THIS EQUATION

Answer by mananth(16949)

2X + 6Y = 0
for x = 3, y=-1 , (3,-1)
for x=6, y=-2(6,-2)
6X - 2Y = 0
for x= 1, y = 3 ( 1,3)
for x=2, y=6 ( 2,6)
Plot the points and find the point of intersection. That is the solution to the system. (0,0)